# Close-in compact super-Earth systems emerging from resonant chains: slow destabilization by unseen remnants of formation

### v1

@@ 3.3 Timing of the instability
The time required for a resonant system to go unstable, and thus transform into a non-resonant system, is of particular importance thanks to recent results of the population characteristics of young systems and an increasing number of clear trends in planetary system architectures with age. Our goal in particular is to determine whether purely N-body dynamics can reproduce the rapid decline in the fraction of systems observed near mean motion resonance on a timescale of approximately (Dai2024). Fig. 8 shows the cumulative distributions for the time of instability of our systems, separated by simulation suite. The “instability” is defined as the first time that a planet with period d at the end of the disk phase had its mass or semi-major axis change by more than 5%, and thus does not include instabilities that occur purely in the outer system that do not propagate to the inner planets.
We found that the instability time distribution is well-modeled by a Weibull distribution (Weibull1951), the cumulative function of which is given by
Here, represents the earliest possible time for an instability, defined to be the disk dissipation time , is a scale parameter, and is a shape parameter. While not derived from a specific physical model, the Weibull distribution is widely used in failure rate statistics to describe the time of failure of a complex system made of multiple components111Indeed, Weibull’s original justification for this distribution was to study the lifetime of a chain that breaks upon the failure of a single link. and its parameters have a straightforward interpretation. The scale parameter is defined simply as the time at which a fraction of systems have failed. Furthermore, in this context, one speaks of the hazard function, which describes the instantaneous rate of failure for systems still functioning at time . For the Weibull distribution, the hazard function takes the form
@@
We fit the simulated instability time distribution of each simulation suite to a Weibull distribution by maximizing the total likelihood. Systems that remained stable over the duration of the integration were assigned likelihood . Uncertainties in the distribution parameters and sample draws from the fit were computed by treating the estimated Hessian matrix returned by the BFGS optimizer implemented in scipy as a covariance matrix (Nocedal2006).
The Weibull parameters best fits and uncertainties are given in Tab. 2. Draws from the fit are also shown in Fig. 8 along with the simulation-derived empirical distribution function. The instability timescale spans several orders of magnitude across simulation suites. There is a clear dependence on total mass that is steeper from to than from to , suggesting that these resonant chains span a stability limit driven by the overlap of resonances (Quillen2011; Petit2020a). The simulations with a pressure bump also have that is a factor of times smaller than the corresponding ones without.
@@ 3.4 Simulated observations
In order to directly compare our results to the observed transiting planet sample, we applied a synthetic observation algorithm to determine which planetary systems would be observed by a transiting survey. To do this, for each simulated system, we simulate observations from 1000 random viewing angles and record which planets transit the star using the CORBITS package (Brakensiek2016), constructing a large synthetic observation catalog.
Our integrations only continued to 500 Myr due to computation time constraints, but typical Kepler systems are much older. We take a fiducial age of . As resonant chains continued going unstable until the end of our simulations, we expect that even fewer will be resonant at late times. We extrapolated the instability fraction to 5 Gyr by sampling the fits to the Weibull distribution and computing the cumulative value at 5 Gyr, taking the median and ranges for each simulation suite. Then, we created a new observed catalog by resampling the previous one according to the new unstable fraction. For example, 57.5% of the 20flat simulations went unstable by 500 Myr, but extrapolating the Weibull fit leads to unstable systems at 5 Gyr. We generate a new observed catalog where 87.6% of systems are drawn from unstable systems and the remainder from the stable systems, while still sampling random viewing angles. Previous work performed similar resampling (Izidoro2017; Izidoro2021a), but only in an ad hoc way to match the observed exoplanet demographics rather than deriving it from the instability statistics.
@@
Figures 9–13 show the inferred multiplicity and period ratio statistics that would be observed by a transiting survey, split into the stable and unstable systems and the resampled catalog in purple. We also plot the multiplicity and period ratio distributions from candidate and confirmed planets detected by the Kepler mission, taken from the Exoplanet Archive.222https://exoplanetarchive.ipac.caltech.edu/ To compare directly with simulations, only planets with periods are included and systems with a giant planet (defined as radius ) are excluded.
@@
As seen in previous studies, planetary systems that underwent a dynamical instability closely resemble the observed transiting planet population. The high mutual inclinations from unstable systems leads to a large number of systems with only a single planet in transit, explaining the so-called Kepler dichotomy between single and multi-transiting systems (Johansen2012). Furthermore, the observed period ratio distribution depends strongly on the total mass, but matches the observed period ratio distribution.
@@
The strength of the produced dichotomy depends strongly on the initial mass. Higher mass planets systems generate larger mutual inclinations during the instability, which directly strengthens the dichotomy (Izidoro2017; Millholland2021). Furthermore, in our simulations, low mass systems have a smaller instability fraction and so the final system catalog is dominated by high multiplicity resonant chains. We find that the 40steep and 40straight simulations most closely reproduce the observed Kepler multiplicity distribution of all of the tested configurations. The lower mass 20flat and 20straight simulations still produce a dichotomy, although it is too weak to match the nearly of single transiting systems observed by Kepler. Finally, the lowest mass 10straight simulation, with its low inclinations and low rate of instability, does not show a distinct dichotomy.
@@
The orbital spacing of planetary systems post-instability is much broader for more massive systems. The observed Kepler period ratio distribution is most closely matched by the unstable systems that started with (orange curve, Figs. 9 and 12). However, the overall period ratio distribution is strongly influenced by the stable systems because, despite their rarity, they are frequently observed as multiples. In the end, a very close match to the observed distribution is obtained in the 40straight run, although 20flat is a satisfactory match.
@@
Now that we have confirmed that the final states of these simulations are an acceptable match to observed mature systems, we consider the new constraint of the time evolution of the fraction of systems near mean-motion resonance. In the absence of robust dynamical information for most systems, Dai2024 defined “near-resonant” systems as having a planet pair with for first-order resonances and for second-order resonances ( is defined in Section 4.2). This period ratio-based definition is different than our stable/unstable distinction because some instabilities can occur without breaking all resonances, and because unstable systems occasionally produce pairs that are serendipitously near MMR in period ratio. For the latter reason, even a catalog of entirely non-resonant systems will be measured as having a 10–20% incidence of resonance by the empirical criterion.
@@
Fig. 14 shows this empirical criterion applied to our five simulation suites as a function of time including the extrapolated instability fractions. The results follow similar trends to Fig. 8. The 10straight and 20straight simulations have too few instabilities to reproduce the near-resonance incidence for systems between 100 Myr and 1 Gyr. On the other hand, 40straight and 40steep have very early instabilities that result in few detected resonant systems at 50 Myr, in conflict with the data. The best result is obtained in the 20flat simulations, which remain mostly resonant at 50 Myr but steadily continue breaking out of resonance through the entire integration. Furthermore, the incidence of resonance in this model extrapolated to 5 Gyr is very close to the measured value.
@@
To summarize briefly, the 10straight and 20straight are a poor fit to the data, being consistently too low mass, too resonant, and too compact. In contrast, 20flat, 40steep, and 40straight qualitatively reproduce many aspects of the data well. However, each of them fails at least one constraint. 20flat has under-excited inclinations, leading to too few single-transiting systems and too compact observed systems. 40steep and 40straight are a better match to the Kepler statistics but have too many early instabilities. It is likely that a mixture of these models would be an even better fit (Goldberg2022). Generically, there is a wide variety of protoplanetary disks and the resulting planetary systems inherit this diversity (Emsenhuber2021a).
@@ 4.1 Why do some systems go unstable?
Fig. 15, showing a typical example of the the 20flat simulation suite, gives a clue as to how these initially resonant systems can slowly unravel. Upon exiting the disk phase, this system contains two resonant chains separated by a large gap of 3.45 in period ratio. The inner chain is relatively widely spaced, with pairs in the 2:1, 3:2, and 4:3 resonances. The outer chain is far more compact, containing pairs in 5:4, 7:6, 8:7, 1:1 (co-orbital), 6:5, and 5:4 resonances. All resonances are librating upon exiting the disk phase, although the outermost pairs have fairly high libration amplitude. From then until , it is clear that pure N-body dynamics are causing the system to chaotically diffuse in action space, increasing the amplitude of libration of some pairs while decreasing that of others. Around , the outermost resonance rapidly grows in libration amplitude and then begins sporadically circulating, which starts to gradually excite the libration of the remaining resonances in the chain. As the eccentricities grow, new resonant modes appear for the co-orbital pair (near ) and the 6:5 resonance enters circulation. At around , the growth in libration amplitudes suddenly accelerates and the entire outer chain breaks simultaneously. Without the protection against close conjunctions provided by the resonant libration, the outer planets excite each other onto crossing orbits and collide, leaving just two survivors.
@@
A particularly remarkable fact is that the inner system appears to be stable on its own. While it receives extra secular precession from the outer system, the separation of timescales between secular and resonant motion ensures that it precesses as a solid object without affecting the mean motion resonances. The outer cluster of planets is unlikely to be observed in transit, especially after gaining significant mutual inclination during the instability, yet plays the key role in the evolution of the entire system. We refer the interested reader to the recent work by Ogihara2025 for details on the importance of secular perturbations from an outer system to instabilities of inner resonant chains.
This two-phase, outer to inner mode is the most common way that resonant chains break in our simulations, but other pathways are possible. Some outer chains go unstable but the inner chain is unaffected (these are counted as stable in our analysis). In a few cases, the instability clearly arose from the inner planets and spread to the outer system. Finally, in other cases it is ambiguous which set of planets is the main source of diffusion. In general, it is not clear why some resonant chains go unstable, although one instability in the three-planet case is understood (Pichierri2020) and heuristics for longer chains are known (Goldberg2022a). We hypothesize that the higher incidence of instability in the simulations with a pressure bump is because low masses and slow migration in the 1 au region consistently creates a very compact resonant chain there that can pass the instability to the inner planets. Planets massive enough to saturate the corotation torques and reach the inner system typically capture into wider resonances which are less prone to resonant overlap and chaos. In contrast, in purely power-law disks (e.g., Fig. 16), there is no clear inner/outer system separation. The outermost planets have very low masses and the entire chain has very low resonant libration amplitude.
@@ 4.2 Remaining resonant systems
A fundamental prediction of the breaking-the-chains model is that the surviving stable systems should closely resemble the observed resonant chains (Bitsch2024). The simplest comparison is to the orbital period architectures; we will focus on period ratios of adjacent planets and proximity to two- and three-body resonances. For each adjacent planet pair with periods , we describe proximity to two-body resonance by the parameter
where we consider only the closest first- or second-order : resonance. To estimate proximity to zeroth-order three-body resonance, we follow the technique of Goldberg2021. For each planet triplet in the system with mean motions , we compute the value
for all positive coprime integer values of and with and normalize it by . Then only the smallest value of is kept for each system.
Fig. 17 shows these statistics for each simulation suite after applying the synthetic observations filter from Section 3.4. We also plot these statistics for the observed 3+ planet resonant chain sample identified by Kajtazi2023 with the addition of TOI-1136 (Dai2023) and again removing systems with giant planets. Simulations suites 20flat and 40steep closely match the observed period ratio distribution, in particular reproducing the strong abundance of 3:2 pairs. In contrast, the simulations with pure power law disks produced too many systems with 5:4 and 4:3 pairs compared to the observed population. However, our simulated planet pairs are much closer to exact resonant commensurability, as expressed by (top right panel), than observed. Nevertheless, the proximity to three-body commensurability is very close the observations, especially for 20flat and 40steep. In other words, many systems show a positive deviation from two-body resonance while being clearly in three-body resonance. The inability of Type I migration to reproduce this feature of the observations is well known (e.g., Charalambous2023, and references therein). Tidal dissipation, which we do not include in our simulations for simplicity, can increase while preserving and potentially explain this mismatch (Papaloizou2015; Goldberg2021; Charalambous2023).
Finally, we note that many of these chains have more detailed mass and eccentricity constraints measured by transit timing variations, but a detailed comparison to these inhomogeneous analyses is left to further work.
@@ 4.3 Neglected effects
Our scheme neglects physics that may be important in planet formation. Real disks are turbulent and these stochastic eddies can suppress capture into resonance during migration (Batygin2017; Chen2025). The true impact will depend on the location of the magneto-rotational instability turn-off (CevallosSoto2025). Nevertheless, this choice appears to be justified. Izidoro2017 considered the effect of stochastic kicks caused by turbulence in their model with a much larger than we use and found that the results were nearly indistinguishable from non-turbulent runs.
Disks are also not static or uniformly depleting in time. Our surface density profile is inspired by the results of Suzuki2016, who consistently found a flat or increasing surface density in the inner disk. However, the position of the peak in surface density evolves outwards in time as the disk dissipates. Ogihara2024 performed N-body simulations with Type I migration directly using these disk models and found similar super-Earth systems to ours. It is likely that the details of the disk profile are most important when planets are and the corotation torque can act as a mass filter. This occurs in our simulations around to , when our disk profiles closely match the results of Suzuki2016.
We also ignored interior structure and atmospheric evolution of the forming planets by always assuming their density to be . In reality, young planets are expected to have puffy, extended envelopes that slowly cool and contract. The connection between this and widespread giant impacts after the gas disk dissipation is not clear. Although SPH simulations of super-Earth and sub-Neptune collisions imply relatively little atmospheric loss from the mechanical shock (Denman2020), Biersteker2019 proposed that post-collision heating would thermally expand the envelope and lead to complete loss via Parker winds. Izidoro2022 showed that this effect, along with expanded radii of water-rich planets that accreted outside of the ice line, can even produce a radius gap between super-Earths and sub-Neptunes.
The radii themselves also have dynamical importance. Large planets are more likely to collide, terminating instabilities before too much dynamical excitation. But it is not clear what the proper radius to use is, because overlap of atmospheres may not be sufficient for collision and merger (see discussion in Kimura2025). We therefore chose as a compromise between core-determined and envelope-determined collision and merger.

### v2

@@ 3.3 Timing of the instability
The time required for a resonant system to go unstable, and thus transform into a non-resonant system, is of particular importance thanks to recent results of the population characteristics of young systems and an increasing number of clear trends in planetary system architectures with age. Our goal in particular is to determine whether purely N-body dynamics can reproduce the rapid decline in the fraction of systems observed near mean motion resonance on a timescale of approximately (Dai et al., 2024). Figure 4 shows the cumulative distributions for the time of instability of our systems, separated by simulation suite. The “instability” is defined as the first time that a planet with period d at the end of the disk phase had its mass or semi-major axis change by more than 5%, and thus does not include instabilities that occur purely in the outer system that do not propagate to the inner planets.
We found that the instability time distribution is well-modeled by a Weibull distribution (Weibull, 1951), the cumulative function of which is given by
Here, represents the earliest possible time for an instability, defined to be the disk dissipation time , is a scale parameter, and is a shape parameter. While not derived from a specific physical model, the Weibull distribution is widely used in failure rate statistics to describe the time of failure of a complex system made of multiple components111Indeed, Weibull’s original motivation for this distribution was to study the lifetime of a chain that breaks upon the failure of a single link. and its parameters have a straightforward interpretation. The scale parameter is defined simply as the time at which a fraction of systems have failed. Furthermore, in this context, one speaks of the hazard function, which describes the instantaneous rate of failure for systems still functioning at time . For the Weibull distribution, the hazard function takes the form
@@
We fit the simulated instability time distribution of each simulation suite to a Weibull distribution by maximizing the total likelihood. Systems that remained stable over the duration of the integration were assigned likelihood . Uncertainties in the distribution parameters and sample draws from the fit were computed by treating the estimated Hessian matrix returned by the BFGS optimizer implemented in scipy as a covariance matrix (Nocedal and Wright, 2006).
The Weibull parameters best fits and uncertainties are given in Table 2. Draws from the fit are also shown in Fig. 4 along with the simulation-derived empirical distribution function. The instability timescale spans several orders of magnitude across simulation suites. There is a clear dependence on total mass that is steeper from to than from to , suggesting that these resonant chains span a stability limit driven by the overlap of resonances (Quillen, 2011; Petit et al., 2020). The simulations with a pressure bump also have that is a factor of times smaller than the corresponding ones without.
@@ 3.4 Simulated observations
In order to directly compare our results to the observed transiting planet sample, we applied a synthetic observation algorithm to determine which planetary systems would be observed by a transiting survey. To do this, for each simulated system, we simulate observations from 1000 random viewing angles and record which planets transit the star using the CORBITS package (Brakensiek and Ragozzine, 2016), constructing a large synthetic observation catalog.
Our integrations only continued to 500 Myr due to computation time constraints, but typical Kepler systems are much older. We take a fiducial age of . As resonant chains continued going unstable until the end of our simulations, we expect that even fewer will be resonant at late times. We extrapolated the instability fraction to 5 Gyr by sampling the fits to the Weibull distribution and computing the cumulative value at 5 Gyr, taking the median and ranges for each simulation suite. Then, we created a new observed catalog by resampling the previous one according to the new unstable fraction. For example, 57.5% of the 20flat simulations went unstable by 500 Myr, but extrapolating the Weibull fit leads to unstable systems at 5 Gyr. We generate a new observed catalog where 87.6% of systems are drawn from unstable systems and the remainder from the stable systems, while still sampling random viewing angles. Previous work performed similar resampling (Izidoro et al., 2017, 2021), but only in an ad hoc way to match the observed exoplanet demographics rather than deriving it from the instability statistics.
@@
Figures 5–9 show the inferred multiplicity and period ratio statistics that would be observed by a transiting survey, split into the stable and unstable systems and the resampled catalog in purple. We also plot the multiplicity and period ratio distributions from candidate and confirmed planets detected by the Kepler mission, taken from the Exoplanet Archive.222https://exoplanetarchive.ipac.caltech.edu/ To compare directly with simulations, only planets with periods are included and systems with a giant planet (defined as radius ) are excluded.
@@
As seen in previous studies, planetary systems that underwent a dynamical instability closely resemble the observed transiting planet population. The high mutual inclinations from unstable systems leads to a large number of systems with only a single planet in transit, explaining the so-called Kepler dichotomy between single and multi-transiting systems (Johansen et al., 2012). Furthermore, the observed period ratio distribution depends strongly on the total mass, but matches the observed period ratio distribution.
@@
The strength of the produced dichotomy depends strongly on the initial mass. Higher mass planets systems generate larger mutual inclinations during the instability, which directly strengthens the dichotomy (Izidoro et al., 2017; Millholland et al., 2021). Furthermore, in our simulations, low mass systems have a smaller instability fraction and so the final system catalog is dominated by high multiplicity resonant chains. We find that the 40steep and 40straight simulations most closely reproduce the observed Kepler multiplicity distribution of all of the tested configurations. The lower mass 20flat and 20straight simulations still produce a dichotomy, although it is too weak to match the nearly of single transiting systems observed by Kepler. Finally, the lowest mass 10straight simulation, with its low inclinations and low rate of instability, does not show a distinct dichotomy.
@@
The orbital spacing of planetary systems post-instability is much broader for more massive systems. The observed Kepler period ratio distribution is most closely matched by the unstable systems that started with (orange curve, Figs. 5 and 8). However, the overall period ratio distribution is strongly influenced by the stable systems because, despite their rarity, they are frequently observed as multiples. In the end, a very close match to the observed distribution is obtained in the 40straight run, although 20flat is a satisfactory match.
@@
Now that we have confirmed that the final states of these simulations are an acceptable match to observed mature systems, we consider the new constraint of the time evolution of the fraction of systems near mean-motion resonance. In the absence of robust dynamical information for most systems, Dai et al. (2024) defined “near-resonant” systems as having a planet pair with for first-order resonances and for second-order resonances ( is defined in Section 4.2). This period ratio-based definition is different than our stable/unstable distinction because some instabilities can occur without breaking all resonances, and because unstable systems occasionally produce pairs that are serendipitously near MMR in period ratio. For the latter reason, even a catalog of entirely non-resonant systems will be measured as having a 10–20% incidence of resonance by the empirical criterion.
@@
Figure 10 shows this empirical criterion applied to our five simulation suites as a function of time including the extrapolated instability fractions. The results follow similar trends to Fig. 4. The 10straight and 20straight simulations have too few instabilities to reproduce the near-resonance incidence for systems between 100 Myr and 1 Gyr. On the other hand, 40straight and 40steep have very early instabilities that result in few detected resonant systems at 50 Myr, in conflict with the data. The best result is obtained in the 20flat simulations, which remain mostly resonant at 50 Myr but steadily continue breaking out of resonance through the entire integration. Furthermore, the incidence of resonance in this model extrapolated to 5 Gyr is very close to the measured value.
@@
To summarize briefly, the 10straight and 20straight are a poor fit to the data, being consistently too low mass, too resonant, and too compact. In contrast, 20flat, 40steep, and 40straight qualitatively reproduce many aspects of the data well. However, each of them fails at least one constraint. 20flat has under-excited inclinations, leading to too few single-transiting systems and too compact observed systems. 40steep and 40straight are a better match to the Kepler statistics but have too many early instabilities. It is likely that a mixture of these models would be an even better fit (Goldberg and Batygin, 2022). Generically, there is a wide variety of protoplanetary disks and the resulting planetary systems inherit this diversity (Emsenhuber et al., 2021b).
@@ 4.1 Why do some systems go unstable?
Figure 11, showing a typical example of the the 20flat simulation suite, gives a clue as to how these initially resonant systems can slowly unravel. Upon exiting the disk phase, this system contains two resonant chains separated by a large gap of 3.45 in period ratio. The inner chain is relatively widely spaced, with pairs in the 2:1, 3:2, and 4:3 resonances. The outer chain is far more compact, containing pairs in 5:4, 7:6, 8:7, 1:1 (co-orbital), 6:5, and 5:4 resonances. All resonances are librating upon exiting the disk phase, although the outermost pairs have fairly high libration amplitude. From then until , it is clear that pure N-body dynamics are causing the system to chaotically diffuse in action space, increasing the amplitude of libration of some pairs while decreasing that of others. Around , the outermost resonance rapidly grows in libration amplitude and then begins sporadically circulating, which starts to gradually excite the libration of the remaining resonances in the chain. As the eccentricities grow, new resonant modes appear for the co-orbital pair (near ) and the 6:5 resonance enters circulation. At around , the growth in libration amplitudes suddenly accelerates and the entire outer chain breaks simultaneously. Without the protection against close conjunctions provided by the resonant libration, the outer planets excite each other onto crossing orbits and collide, leaving just two survivors.
@@
A particularly remarkable fact is that the inner system appears to be stable on its own. While it receives extra secular precession from the outer system, the separation of timescales between secular and resonant motion ensures that it precesses as a solid object without affecting the mean motion resonances (Petit et al., 2017; Tamayo et al., 2021). The outer cluster of planets is unlikely to be observed in transit, especially after gaining significant mutual inclination during the instability, yet plays the key role in the evolution of the entire system. We refer the interested reader to the recent work by Ogihara and Kunitomo (2026) for details on the importance of secular perturbations from an outer system to instabilities of inner resonant chains.
This two-phase, outer to inner mode is the most common way that resonant chains break in our simulations, but other pathways are possible. Some outer chains go unstable but the inner chain is unaffected (these are counted as stable in our analysis). In a few cases, the instability clearly arose from the inner planets and spread to the outer system. Finally, in other cases it is ambiguous which set of planets is the main source of diffusion.
In general, it is not clear why some resonant chains go unstable, although one instability in the three-planet case is understood (Pichierri and Morbidelli, 2020) and heuristics for longer chains are known (Goldberg et al., 2022). We hypothesize that the higher incidence of instability in the simulations with a pressure bump is because low masses and slow migration in the 1 au region consistently create a very compact resonant chain. Although this outer chain is stable to chaotic overlap of adjacent first-order resonances (Deck et al., 2013), a large network of higher-order resonances can drive diffusion away from the resonant center and, once the instability is eventually triggered, an impulsive kick can be passed to the inner planets. Instabilities rarely occur in the inner chain because planets massive enough to saturate the corotation torques and reach the inner system typically capture into wider resonances which are less prone to resonant overlap and chaos. In contrast, in purely power-law disks (e.g., Fig. 12), there is no clear inner/outer system separation. The outermost planets have very low masses and the entire chain has very low resonant libration amplitude.
@@ 4.2 Remaining resonant systems
A fundamental prediction of the breaking-the-chains model is that the surviving stable systems should closely resemble the observed resonant chains (Bitsch and Izidoro, 2024). The simplest comparison is to the orbital period architectures; we will focus on period ratios of adjacent planets and proximity to two- and three-body resonances. For each adjacent planet pair with periods , we describe proximity to two-body resonance by the parameter
where we consider only the closest first- or second-order : resonance. To estimate proximity to zeroth-order three-body resonance, we follow the technique of Goldberg and Batygin (2021). For each planet triplet in the system with mean motions , we compute the value
for all positive coprime integer values of and with and normalize it by . Then, only the smallest value of is kept for each system.
Figure 13 shows these statistics for each simulation suite after applying the synthetic observations filter from Section 3.4. We also plot these statistics for the observed 3+ planet resonant chain sample identified by Kajtazi et al. (2023) with the addition of TOI-1136 (Dai et al., 2023) and again removing systems with giant planets. Simulations suites 20flat and 40steep closely match the observed period ratio distribution, in particular reproducing the strong abundance of 3:2 pairs. In contrast, the simulations with pure power law disks produced too many systems with 5:4 and 4:3 pairs compared to the observed population. However, our simulated planet pairs are much closer to exact resonant commensurability, as expressed by (top right panel), than observed. Nevertheless, the proximity to three-body commensurability is very close the observations, especially for 20flat and 40steep. In other words, many systems show a positive deviation from two-body resonance while being clearly in three-body resonance. The inability of Type I migration to reproduce this feature of the observations is well known (e.g., Charalambous et al., 2023, and references therein). Tidal dissipation, which we do not include in our simulations for simplicity, can increase while preserving and potentially explain this mismatch (Papaloizou, 2015; Goldberg and Batygin, 2021; Charalambous et al., 2023).
Finally, we note that many of these chains have more detailed mass and eccentricity constraints measured by transit timing variations (TTVs), but a detailed comparison to these inhomogeneous analyses is left to further work.
@@ 4.3 Neglected effects
Our scheme neglects physics that may be important in planet formation. Real disks are turbulent and these stochastic eddies can suppress capture into resonance during migration (Batygin and Adams, 2017; Chen et al., 2025). The true impact will depend on the location of the magneto-rotational instability turn-off (Cevallos Soto and Zhu, 2025). Nevertheless, this choice appears to be justified. Izidoro et al. (2017) considered the effect of stochastic kicks caused by turbulence in their model with a much larger than we use and found that the results were nearly indistinguishable from non-turbulent runs.
Disks are also not static or uniformly depleting in time. Our surface density profile is inspired by the results of Suzuki et al. (2016), who consistently found a flat or increasing surface density in the inner disk. However, the position of the peak in surface density evolves outwards in time as the disk dissipates. Ogihara et al. (2024) performed N-body simulations with Type I migration directly using these disk models and found similar super-Earth systems to ours. It is likely that the details of the disk profile are most important when planets are and the corotation torque can act as a mass filter. This occurs in our simulations around to , when our disk profiles closely match the results of Suzuki et al. (2016).
We also ignored interior structure and atmospheric evolution of the forming planets by always assuming their density to be . In reality, young planets are expected to have puffy, extended envelopes that slowly cool and contract. The connection between this and widespread giant impacts after the gas disk dissipation is not clear. Although SPH simulations of super-Earth and sub-Neptune collisions imply relatively little atmospheric loss from the mechanical shock (Denman et al., 2020), Biersteker and Schlichting (2019) proposed that post-collision heating would thermally expand the envelope and lead to complete loss via Parker winds. Izidoro et al. (2022) showed that this effect, along with expanded radii of water-rich planets that accreted outside of the ice line, can even produce a radius gap between super-Earths and sub-Neptunes.
The radii themselves also have dynamical importance. Large planets are more likely to collide, terminating instabilities before too much dynamical excitation. But, it is not clear what the proper radius to use is, because overlap of atmospheres may not be sufficient for collision and merger (see discussion in Kimura et al., 2025). We therefore chose as a compromise between core-determined and envelope-determined collision and merger.